Joana’s monthly income is $5000. She spends 40% of her income on grocery and the rest on buying designer clothes. The Provincial Government thinks it is unfair that people spend more than 35% of their income for grocery. In order to lower the proportion of income going to grocery, the Provincial Government gives Joana $500. Given that Joana’s income elasticity of grocery is 2, does the Provincial Government accomplished its goal? Support your answer.
1
Expert's answer
2020-04-08T09:51:38-0400

Joana’s monthly income new"=5000+500=" $"5500;"

Edi "=\frac {\Delta Qx}{\Delta I}×\frac {I}{Qx};"

"\Delta Qx=\frac {(5500-5000)}{5000}×100=" %"10;"

"I=40\u00d75000\/100=" $ "2000;"

"\frac {10}{\Delta I}×\frac {2000}{5000}=2;"

"\frac {10}{\Delta I}×0.4=2;"

"\frac {10}{\Delta x}=5;"

"\Delta I=" %"2;"

"I"new"=2000+\frac {2×2000}{100};"

"I" new"=2040;"

% of her income on grocery"=2040\/5500\u00d7100=" %"37;"

answer: the government did not achieve its goal, as Joanna will spend on products 37% of its income, rather than 35%.